OpenStudy (anonymous):

You are going on a business trip for a manufacturing company to Italy and need to figure out the currency conversion. You are going to take 1500 US dollars. After checking the internet you find the exchange rate is 1 US dollar equals 0.74 Euros. a. Write the equation to convert US dollars to Euros. b. How many Euros will you have if you convert all of the 1500 US dollars? c. On your return trip you need to calculate how much US dollars you will have left over. To do this you need to find the inverse of your previous equation. What is the inverse equation to calculate how much US dollar

OpenStudy (anonymous):

you will have? d. After your trip you have 75 Euros left. How much US dollars will you have? The Italian company uses temperatures in Celsius and Kelvin for the different manufacturing processes. However, in America, the company uses Fahrenheit. The formula for converting Celsius to Kelvin is represented by k(x)below and the formula for converting Fahrenheit to Celsius is represented by c(x) below: k(x) = x + 273.16 c(x) = 5/9(x - 32) e. For the American companies we need to find the Celsius to Fahrenheit formula by using the inverse. What is this formula? f. Also for the Italian company we need to use the composition k(c(x)) to convert Fahrenheit to Kelvin. What is this formula?

OpenStudy (xapproachesinfinity):

what do you think about the first question

OpenStudy (xapproachesinfinity):

can you figure out that equation

OpenStudy (xapproachesinfinity):

we are looking here for euro so that would be our output yes let's call it E

OpenStudy (xapproachesinfinity):

now there's one other variable we want here, which is the amount of dollar we want to transfer let's call it d for dollar

OpenStudy (xapproachesinfinity):

the relation here is 1 us dollar = 0.74 euro we want to have an equation that will give the conversion for any d amount

OpenStudy (xapproachesinfinity):

E=0.74d is the equation

OpenStudy (anonymous):

i got a. 1/0.74 = 1500/x b. 1110euros c. 1/0.74 = x d. c(x) = 5/9(x - 32) just say x = 5/9(y - 32) and solve for y e.Multiply by 9, then divide by 5, then add 32 f would be <SUBSTITUTE c(x) here> + 273.16 which would be (5/9(x - 32) ) + 273.16